OpenStudy (marziajackson$):

What is the value of the expression? 34 + 7 *6^2

OpenStudy (nincompoop):

do you know how to multiply and add?

OpenStudy (marziajackson$):

yes

Nnesha (nnesha):

PEMDAS! Parentheses Exponent Mulitply Divide Add Subtract

OpenStudy (nincompoop):

okay then: \(6^2 \) means \(6 \times 6 \) so you have the following 34 + 7*6*6 then apply the ORDER of operation multiplication then addition

OpenStudy (mikeymaximum):

I got 286...

OpenStudy (mikeymaximum):

6^2=36

OpenStudy (mikeymaximum):

7*36=252

OpenStudy (mikeymaximum):

252+34=286

Nnesha (nnesha):

okay .... now let her solve this let's what she get

OpenStudy (marziajackson$):

I see it now. The answer Is 286

Nnesha (nnesha):

are you sure ??

OpenStudy (marziajackson$):

yeah that's what I got the second time

Nnesha (nnesha):

well................. 286 is right first multiply and then add 7 times 6^2 and then add 34 .........